OpenStudy (anonymous):

factor completely: 4y²+20y+25 Please explain how you did it! Thank you!

OpenStudy (anonymous):

you are supposed to recognize this as a "perfect square" \[4y^2\] is the square of \[2y\],25 is the square of 5, and \[20y=2\times 2y\times 5\]

OpenStudy (anonymous):

so it factors as \[(2y+5)^2\] the explanation came first, then the answer

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

Yeah were supposed to make it a perfect square

OpenStudy (anonymous):

But I don't know how to do a perfect square with three terms.

OpenStudy (anonymous):

it looks like this: \[a^2+2ab+b^2=(a+b)^2\]

OpenStudy (anonymous):

in your case \[a=2y, b=5, 2ab=20y\]

OpenStudy (anonymous):

So 20y would stay the same?

OpenStudy (anonymous):

so is this the answer: 2y+20y+5=(2y+5)²

OpenStudy (anonymous):

no. you started with \[4y^2+20y+25\] and that is equal to \[(2y+5)^2\]

OpenStudy (anonymous):

I'm really confused.